# Maximum entropy models of neuronal populations at and off criticality

### v1

@@ IV.1 Modulation of ME modeling parameters across network states
A similar behavior is found in our neural network model. In the K-pairwise ME modeling, one constrains the distribution $P(K)$, which provides, at most, an additional NN potentials, $V_{K}$. When there is little or no synchronous firing across the network, as in the subcritical state, $V_{K}$ are mostly zero. On the contrary, $V_{K}$ are not negligible in critical and supercritical states. Notably, we found that $V_{K}$ are strongly altered in the transition from criticality to supercriticality, in particular at large $K$, where they become strongly positive—in the same range $V_{K}$ are zero at criticality. Among all the inferred parameters, the fields $V_{K}$, show the most striking difference when comparing critical and supercritical states. Because they are related to a collective variable, the modulations in $V_{K}$ can be easily translated into changes in avalanche dynamics, as outlined above.
@@ Acknowledgements.
L.d.A. acknowledges support from the Italian MUR project PRIN2017WZFTZP and from NEXTGENERATIONEU (NGEU) funded by the Ministry of University and Research (MUR), National Recovery and Resilience Plan (NRRP), and project MNESYS (PE0000006)-A multiscale integrated approach to the study of the nervous system in health and disease (DN. 1553 11.10.2022). H.J.H. thanks the INCT NeuroComp and the Funcap for support. FL acknowledges support from the European Union’s Horizon research and innovation program under the Marie Sklodowska-Curie Grant Agreement No. 101066790 and from the program TAlent in ReSearch@University of Padua – STARS@UNIPD (project BRAINCIP—Brain criticality and information processing).

### v2

@@ IV.1 Modulation of ME modeling parameters across network states
A similar behavior is found in our neural network model. In the K-pairwise ME modeling, one constrains the distribution $P(K)$, which provides, at most, an additional $N$ potentials, $V_{K}$. When there is little or no synchronous firing across the network, as in the subcritical state, $V_{K}$ are mostly zero. On the contrary, $V_{K}$ are not negligible in critical and supercritical states. Notably, we found that $V_{K}$ are strongly altered in the transition from criticality to supercriticality, in particular at large $K$, where they become strongly positive—in the same range $V_{K}$ are zero at criticality. Among all the inferred parameters, the fields $V_{K}$, show the most striking difference when comparing critical and supercritical states. Because they are related to a collective variable, the modulations in $V_{K}$ can be easily translated into changes in avalanche dynamics, as outlined above.
@@ Acknowledgements.
L.d.A. acknowledges support from the Italian MUR project PRIN2017WZFTZP and from NEXTGENERATIONEU (NGEU) funded by the Ministry of University and Research (MUR), National Recovery and Resilience Plan (NRRP), and project MNESYS (PE0000006)-A multiscale integrated approach to the study of the nervous system in health and disease (DN. 1553 11.10.2022). H.J.H. thanks the INCT NeuroComp and the Funcap for support. FL acknowledges support from the European Union’s Horizon research and innovation program under the Marie Sklodowska-Curie Grant Agreement No. 101066790 and from the program TAlent in ReSearch@University of Padua – STARS@UNIPD (project BRAINCIP—Brain criticality and information processing). This research was supported by the Intramural Research Program of the National Institutes of Mental Health (NIMH), USA, ZIAMH002797, ZIAMH002971.